# Vosh: Voxel-Mesh Hybrid Representation for Real-Time View Synthesis

### v1

@@ Abstract.
The neural radiance field (NeRF) has emerged as a prominent methodology for synthesizing realistic images of novel views. While neural radiance representations based on voxels or mesh individually offer distinct advantages, excelling in either rendering quality or speed, each has limitations in the other aspect. In response, we propose a pioneering hybrid representation named Vosh, seamlessly combining both voxel and mesh components in hybrid rendering for view synthesis. Vosh is meticulously crafted by optimizing the voxel grid of NeRF, strategically with selected voxels replaced by mesh. Therefore, it excels in fast rendering scenes with simple geometry and textures through its mesh component, while simultaneously enabling high-quality rendering in intricate regions by leveraging voxel component. The flexibility of Vosh is showcased through the ability to adjust hybrid ratios, providing users the ability to control the balance between rendering quality and speed based on flexible usage. Experimental results demonstrates that our method achieves commendable trade-off between rendering quality and speed, and notably has real-time performance on mobile devices.
@@
Neural Radiance Field, View Synthesis, Real-Time Rendering.
journal: TOGcopyright: noneccs: Computing methodologies Reconstructionccs: Computing methodologies Neural networksccs: Computing methodologies Volumetric models
@@
Figure 1. Left: The hybrid representation Vosh, combines both voxel component (highlighted with pseudo colors) and mesh component (indicated by gray triangles) in hybrid rendering. Right: The proposed method facilitates real-time view synthesis on various mobile devices, e.g., laptops and mobile phones.
1. Introduction
@@
The novel view synthesis based on the Neural Radiance Field (NeRF) leverages neural networks to model three-dimensional scenes and generates images from specified viewpoints (Mildenhall et al., 2021). In general, NeRF-based approaches demonstrate proficiency in producing high-quality rendering for novel views with sparse inputs, effectively handling materials with specular and translucent properties. This versatility positions them for diverse applications in autonomous navigation (Kwon et al., 2023), virtual reality/augmented reality (Li et al., 2022a), and beyond (Sun et al., 2022b; Jiang et al., 2023). Despite these merits, current methods face challenges in achieving a balance between high-quality rendering and real-time rendering, particularly when deployed on mobile devices.
@@
The rendering speed of neural radiance fields is significantly influenced by the chosen representation, with notable implications for sampling requirements and query costs (Yu et al., 2021). Typically, the original NeRF uses deep multilayer perceptrons (MLPs) to model neural radiance fields, relying on dense sampling for querying to approximate integrals in volume rendering. Unfortunately, this approach typically causes excessively long rendering times. A substantial amount of effort has explored alternatives such as implicit voxel representations (Sun et al., 2022a; Hedman et al., 2021; Liu et al., 2020; Fridovich-Keil et al., 2022; Müller et al., 2022; Reiser et al., 2021) and explicit mesh representations (Chen et al., 2023; Tang et al., 2023; Rakotosaona et al., 2023; Yariv et al., 2023), aiming to exclude irrelevant regions for computation. These methods reduce the number of sampling points and employ smaller MLPs to expedite rendering. However, even though approaches based on implicit voxels can achieve high-quality rendering in complex scenes, the dense sampling in non-empty regions is prohibitively expensive.
@@
Methods based on mesh representation have demonstrated significantly accelerated rendering speeds, enabling real-time rendering on common mobile devices. However, the challenge lies in the reconstruction of scene-level mesh structure, particularly in areas with intricate geometry and textures (see Figure 1), such as thin structures or undetermined regions like backgrounds (Reiser et al., 2023). It is observed that meshes can match the rendering quality akin to voxels in regions with simple geometry and textures, as well as offering highly efficient rendering. Therefore, a promising solution involves employing hybrid representations for reconstruction towards both high-quality rendering and fast speed. Existing works on hybrid representations (Wang et al., 2023; Guo et al., 2023), focus primarily on object or proximal region reconstruction, thus limiting flexibility to represent complex scenes and control over different parts of the hybrid representation. Alternatively, we leverage the strengths of voxels and mesh, to design a hybrid representation towards a better balance between quality and speed, especially enabling high-quality rendering on mobile devices (see Figure 1).
@@
In summary, the main contribution of this paper is a new hybrid representation namely Vosh (VOxel-meSH), by combing voxels and mesh for NeRF-based real-time view synthesis. It also allows for a dynamic control of the balance between voxel and mesh components by voxel adjustment. Such flexibility results in a versatile spectrum of NeRF representations, catering to the varied computational capabilities that facilitate real-time view synthesis across different devices. Experimental results exhibit the super performance of our method compared to state-of-the-art methods.
Figure 2. An overview of the proposed methodology. The training phase starts from grid training for obtaining initial voxels. Then, a portion of voxels becomes mesh in the voxels to mesh conversion. Subsequently, the combination of voxels and mesh are optimized through hybrid rendering and voxels adjustment to obtain the final hybrid representation Vosh. The inference phase realizes real-time hybrid rendering with Vosh even on mobile phones.
2. Related Work
@@
There are numerous published approaches to view synthesis and NeRF, and we refer interested readers to (Tewari et al., 2022; Gao et al., 2022) for a comprehensive overview. In this context, we focus on the real-time view synthesis and NeRF representations most related to our method.
2.1. Real-time view synthesis
@@
NeRF (Mildenhall et al., 2021) as a well-known technology for view synthesis, has been the focus of numerous endeavors aimed at enhancing its rendering speed while maintaining high-quality rendering. SNeRG (Hedman et al., 2021) achieves fast rendering without CUDA, using sparse voxel representations and a compact neural network for deferred rendering. Plenoxels (Fridovich-Keil et al., 2022) and PlenOctrees (Yu et al., 2021) expedite ray marching with sparse voxel representation, refining voxels near object surfaces. MERF (Reiser et al., 2023) enhances rendering speed and spatial occupancy by replacing pure voxel grids with a low-resolution voxel grid and a high-resolution triplane projection representation (Chan et al., 2022). These methods mostly excel on high-end GPUs, whereas real-time rendering on mobile devices still remains a challenge.
@@
MobileNeRF (Chen et al., 2023) achieves fast rendering by constraining NeRF to triangles and using rasterization. NeRF2Mesh (Tang et al., 2023) uses NeRF to initialize geometric structures, iteratively refining mesh attributes to improve rendering quality. BakedSDF (Yariv et al., 2023) optimizes a blended neural signed distance function representation, enabling real-time rendering of complex scenes on a high-quality mesh. While mesh-based methods offer fast rendering, they may suffer quality loss compared to voxel-based methods.
@@
3D Gaussian Splatting (Kerbl et al., 2023) creates a 3D Gaussian representation around each point starting from a point cloud model, utilizing the Splatting method for real-time rendering of high-resolution scenes. But it suffers elongated and popping artifacts, and high memory usage for deployment on mobile phones. In contrast, our hybrid representation combines the strengths of voxel and mesh representations, allowing high-quality rendering of complex scenes on mobile devices.
2.2. Hybrid Scene Representation
@@
Combining explicit mesh representation with implicit fields introduces functionalities such as NeRF editing and rendering acceleration. NeuMesh (Yang et al., 2022) encodes a neural implicit field with disentangled geometry and texture on mesh vertices, enabling a range of neural rendering-based editing operations. EyeNeRF (Li et al., 2022b) integrates an explicit surface with an implicit volume representation to model the human eye. Mesh-Aware-RF (Qiao et al., 2023) embeds a known polygonal mesh into NeRF, employing ray tracing and ray marching alternately to update radiance along rays.
@@
The hybrid representation like Adaptive Shells (Wang et al., 2023) and VMesh (Guo et al., 2023) has been used in view synthesis. Adaptive Shells confines volume rendering to a narrow band around the object for effective NeRF rendering. VMesh, closely related to our work, uses a textured mesh and auxiliary sparse voxels but differs by employing SDF field and volume density field to simultaneously learn the entire scene. However, VMesh relies on gradient descent optimization that lacks controllability and flexibility. Indeed, VMesh mainly focuses on rendering individual objects or bounded scenes, while the proposed method in this paper can handle more scenes, such as unbounded outdoor scenes.
3. Method Overview
@@
Concretely, the training phase consists of three stages: grid training, voxels to mesh conversion and optimization. In the grid training stage, we train a voxel grid model through volume rendering. Then, leveraging the trained voxel grid, we extract an explicit surface and train the surface using differentiable surface rendering for both appearance and geometry optimization. This converts the initial voxel grid into a combination of voxel and mesh components. Finally, the voxels and mesh are further optimized through a training process involving hybrid rendering and voxels adjustment to obtain the desired hybrid representation. Once training is complete, we export assets containing sparse voxels and textured meshes. Subsequently, real-time view synthesis on mobile devices is implemented using WebGL. Next, we elaborate the algorithm details in the training and inference phases.
4. Hybrid Representation construction
The construction of a hybrid representation in the training phase commences with the initialization of a volume comprising solely of voxels. This volume is then subject to a conversion process wherein a portion of the voxels is replaced with an explicit mesh according to rendering quality on par with the original voxel representation.
4.1. Grid training for initial voxels
@@
Initially, the entire scene is reconstructed using a high-resolution voxel grid. It retains sufficient geometric and appearance information essential for subsequent mesh extraction and hybrid representation. Here, we employ SNeRG++ (an improved version of SNeRG (Hedman et al., 2021) in MERF (Reiser et al., 2023)) as the benchmark for our voxel grid, due to its capability to effectively model unbounded scenes and generate real-time rendering assets.
@@
Concretely, SNeRG++ is built upon the volume rendering principles of vanilla NeRF, which involves sampling along the direction of a ray originating from the ray origin $o$ along the direction $d$ to obtain sample points $\mathrm{\mathbf{x}}_{i}=o+t_{i}\mathrm{\mathbf{d}}$. Based on the spatial position of these sample points, it uses the neural radiance field to calculate volumetric density $\sigma_{i}$, resulting in rendering weights $w_{i}$:
(1) $w_{i}=\alpha_{i}T_{i},\;\;\;\;\;\;\;\;T_{i}=\prod^{i-1}_{j=1}(1-\alpha_{j}),\;\;\;\;\;\;\;\;\alpha_{i}=1-e^{-\sigma_{i}\delta_{i}},$
@@
where $T_{i}$, $\alpha_{i}$ and $\delta_{i}$ represent the transmittance, opacity and spacing of sampling point $i$, respectively. SNeRG++ adopts the concept of deferred rendering, decomposing the appearance information of the neural radiance field into diffuse color $\mathrm{\mathbf{c}}_{d,i}$ and view-dependent features $f_{i}$. By cumulatively sampling the appearance information along the ray, as shown in Figure 4a, we obtain the diffuse color $C_{d}$ and view-dependent features $F$ of the ray as
(2) $\mathrm{\mathbf{C}}_{d}=\sum_{i}w_{i}\mathrm{\mathbf{c}}_{d,i},\;\;\;\;\;\;\;\;\mathrm{\mathbf{F}}=\sum_{i}w_{i}\mathrm{\mathbf{f}}_{i}.$
@@
Finally, the view-dependent color is calculated by inputting the diffuse color $\mathrm{\mathbf{C}}_{d}$, view-dependent features $F$, and the ray direction $d$ into a miniature MLP. The result is then added to the diffuse color $\mathrm{\mathbf{C}}_{d}$ to obtain the final color $C$ of the ray.
@@
Besides, in order to enhance the modeling of regions situated far from the center within unbounded scenes and facilitate efficient ray marching calculations, SNeRG++ incorporates a contraction function founded on piecewise projection. This contraction function is defined as follows:
[…]
where $||\cdot||_{\infty}$ denotes the infinity norm, defined as $||x||_{\infty}=\text{max}_{j}|x_{j}|$.
4.2. Voxels to mesh conversion

### v2

# Voxel-Mesh Hybrid Representation For Real-Time View Synthesis By Meshing Density Field

@@ Abstract
The neural radiance fields (NeRF) have emerged as a prominent methodology for synthesizing realistic images of novel views. While neural radiance representations based on voxels or mesh individually offer distinct advantages, excelling in either rendering quality or speed, each has limitations in the other aspect. In response, we propose a hybrid representation named Vosh, seamlessly combining both voxel and mesh components in hybrid rendering for view synthesis. Vosh is meticulously crafted by optimizing the voxel grid based on neural rendering, strategically meshing a portion of the volumetric density field to surface. Therefore, it excels in fast rendering scenes with simple geometry and textures through its mesh component, while simultaneously enabling high-quality rendering in intricate regions by leveraging voxel component. The flexibility of Vosh is showcased through the ability to adjust hybrid ratios, providing users the ability to control the balance between rendering quality and speed based on flexible usage. Experimental results demonstrate that our method achieves commendable trade-off between rendering quality and speed, and notably has real-time performance on mobile devices. The interactive web demo and code are available at https://zyyzyy06.github.io/Vosh.
Index Terms:
@@
Neural Radiance Fields, View Synthesis, Real-Time Rendering.
@@
Figure 1: Left: The hybrid representation Vosh, combines both voxel component (highlighted with pseudo colors) and mesh component (indicated by gray triangles) in hybrid rendering. Right: The proposed method facilitates real-time view synthesis on various mobile devices, e.g., laptops and mobile phones.
I Introduction
@@
The novel view synthesis based on the Neural Radiance Fields (NeRF) leverages neural networks to model three-dimensional scenes and generates images from specified viewpoints [1]. In general, NeRF-based approaches demonstrate proficiency in producing high-quality rendering for novel views with sparse inputs, effectively handling materials with specular and translucent properties. This versatility positions them for diverse applications in autonomous navigation [2], virtual reality/augmented reality [3], and beyond [4, 5]. Despite these merits, current methods face challenges in achieving a balance between high-quality rendering and real-time rendering, particularly when deployed on mobile devices.
@@
The rendering speed of neural radiance fields is significantly influenced by the chosen representation, with notable implications for sampling requirements and query costs [6]. Typically, the original NeRF uses deep multilayer perceptrons (MLPs) to model neural radiance fields, relying on dense sampling for querying to approximate integrals in volume rendering. Unfortunately, this approach typically causes excessively long rendering times. A substantial amount of effort has explored alternatives such as voxel representations [7, 8, 9, 10, 11, 12] and mesh representations [13, 14, 15, 16], aiming to exclude irrelevant regions for computation. These methods reduce the number of sampling points and employ smaller MLPs or no MLPs to expedite rendering. However, even though approaches based on voxels can achieve high-quality rendering in complex scenes, the dense sampling in non-empty regions is prohibitively expensive. 3DGS [17], with its Gaussian representation and Splatting rendering method, also reduces the number of sampled points in rays, achieving real-time rendering of high-resolution scenes. But its high memory usage and platform-related sorting strategies hinder its widespread applications, especially on mobile devices.
@@
The methods based on mesh representation have demonstrated significantly accelerated rendering speeds, enabling real-time rendering on common mobile devices. However, the challenge lies in the reconstruction of scene-level mesh structure, particularly in areas with intricate geometry and textures (see Figure 1), such as thin structures or undetermined regions like backgrounds [18]. It is observed that meshes can match the rendering quality akin to voxels in regions with simple geometry and textures, as well as offering highly efficient rendering. Therefore, a promising solution involves employing hybrid representations for reconstruction towards both high-quality rendering and fast speed.
Existing works on hybrid representations [19, 20], focus primarily on object or proximal region reconstruction, thus abandoning the reconstruction ability of volume representation for complex scenes, such as background regions lacking observation in unbounded outdoor environments. VMesh [20] mainly relies on the signed distance function for the scene reconstruction. However, it may result in low-quality reconstruction and is space-consuming for complex background areas of unbounded scenes. We utilize implicit volumetric density field, which can recover more comprehensive information and be applied in both bounded and unbounded scenes. By partially meshing the density field, a more expressive mesh representation can be obtained. Then we leverage the strengths of voxels and mesh, to design a hybrid representation towards a better balance between quality and speed, especially enabling high-quality rendering on mobile devices (see Figure 1).
@@
In summary, the main contribution of this paper is a new hybrid representation namely Vosh (VOxel-meSH), by combing voxels and mesh for NeRF-based real-time view synthesis. It also allows for a dynamic control of the balance between voxel and mesh components. Such flexibility results in a versatile spectrum of NeRF representations, catering to the varied computational capabilities that facilitate real-time view synthesis across different devices. Experimental results exhibit the superior performance of our method compared to state-of-the-art methods.
II Related Work
@@
There are numerous published approaches to view synthesis and NeRF, and we refer interested readers to [21, 22, 23] for a comprehensive overview. In this context, we focus on the real-time view synthesis and NeRF representations most related to our method.
II-A Real-time view synthesis
@@
NeRF [1] as a well-known technology for view synthesis, has been the focus of numerous endeavors aimed at accelerating its rendering speed while maintaining high quality of rendered images. SNeRG [8] achieves fast rendering without CUDA, using sparse voxel representations and a compact neural network for deferred rendering. Plenoxels [10] and PlenOctrees [6] expedite ray marching with sparse voxel representation, refining voxels near object surfaces. MERF [18] enhances rendering speed and spatial occupancy by replacing pure voxel grids with a low-resolution voxel grid and a high-resolution triplane projection representation [24]. These methods mostly excel on high-end GPUs, whereas real-time rendering on mobile devices still remains a challenge.
@@
MobileNeRF [13] achieves fast rendering by constraining NeRF to triangles and using rasterization. NeRF2Mesh [14] uses NeRF to initialize geometric structures, iteratively refining mesh attributes to improve rendering quality. BakedSDF [16] optimizes a blended neural signed distance function representation based on VolSDF [25], enabling real-time rendering of complex scenes on a high-quality mesh. In order to represent complex backgrounds, BakedSDF uses a contracted space containing Eikonal loss which might have negative impact on the quality of the foreground as claimed in [26]. While mesh-based methods offer fast rendering, they suffer quality loss compared to volume-based methods, especially for unbounded scenes with complex backgrounds.
@@
3D Gaussian Splatting [17] creates a 3D Gaussian representation around each point starting from a point cloud model, utilizing the Splatting method for real-time rendering of high-resolution scenes. To achieve real-time training and rendering, 3DGS relies on CUDA-specific APIs. So original 3DGS is difficult to render on most mobile devices. While recent re-implementations seem to render 3DGS models on commodity hardware, they rely on approximations to sort order and view-dependency whose impact on quality and speed has not yet been evaluated [27]. In contrast, our hybrid representation combines the strengths of voxel and mesh representations, allowing high-quality rendering of complex scenes on mobile devices.
Recently, 3DGS-based compression methods have emerged to effectively reduce the number of Gaussians. Eagles [28] proposes to use quantized embeddings and a coarse-to-fine training strategy to significantly reduce memory storage overhead. F-3DGS [29] represents dense clusters of Gaussians with significantly fewer Gaussians through efficient factorization. While these methods have successfully enhanced 3DGS in terms of storage efficiency and rendering speed, they still require as much or even more VRAM in the rendering compared to the original 3DGS as reported in these papers, which are still not suitable for the use on mobile devices.
II-B Hybrid scene representation
@@
Combining explicit mesh representation with implicit fields paves a new way to NeRF-based processing, such as NeRF editing and rendering acceleration. NeuMesh [30] encodes a neural implicit field with disentangled geometry and texture on mesh vertices, enabling a range of neural rendering-based editing operations. EyeNeRF [31] integrates an explicit surface with an implicit volume representation to model the human eye. Mesh-Aware-RF [32] embeds a known polygonal mesh into NeRF, employing ray tracing and ray marching alternately to update radiance along rays.
@@
The hybrid representation like Adaptive Shells [19] and VMesh [20] has been used in view synthesis. Adaptive Shells confine volume rendering to a narrow band around the object for effective NeRF rendering. VMesh, closely related to our work, uses a textured mesh and auxiliary sparse voxels but differs by employing the signed distance function (SDF) and volume density field to simultaneously learn the bounded scenes or objects. Both methods use NeuS [33] as the implementation for SDF. NeuS learns neural implicit surfaces through volume rendering. For the unbounded scenes with backgrounds, NeuS needs to use NeRF for background modeling that is difficult to represent by itself, which is also a hybrid representation.
It should be noted that VMesh most closely resembling our method in this paper, relies on SDF with density field to capture scene information, while our method only uses the density field. Specifically, when using SDF to reconstruct complex areas like cluttered backgrounds in unbounded scenes, it might lead to low-quality reconstruction and significant space occupation [16]. Due to this limitation, VMesh is primarily suited for single objects or bounded scenes, whereas our density-field-based method effectively manages a wide range of scenarios, including unbounded outdoor scenes. Besides, VMesh requires manual tuning of multiple loss weights for each scene to achieve optimal surface-volume separation as indicated in the paper [20]. Without these adjustments, only a single representation may be optimized. In contrast, our method based on the density field does not rely on scene-specific hyper-parameters, enabling the high-quality hybrid representation for various scenes.
Figure 2: An overview of the proposed methodology. The training phase starts from grid training for obtaining initial voxels. Then, a portion of initial volumetric density field are meshed. Subsequently, the combination of voxels and mesh are optimized through hybrid rendering and voxel pruning to obtain the final hybrid representation Vosh. The inference phase realizes real-time hybrid rendering with Vosh even on mobile phones.
III Method Overview
@@
Concretely, the training phase consists of three stages: grid training, meshing density field and optimization. In the grid training stage, we train a voxel grid model through volume rendering. Then, we determine the region of meshing in the density field based on the difference in rendering quality between the voxel grid and the potential surfaces it contains. This converts the initial voxel grid into a combination of voxel and mesh components. Finally, the voxels and mesh are further optimized through a training process involving hybrid rendering and voxel pruning to obtain the desired hybrid representation. Once training is complete, we export assets containing sparse voxels and textured meshes. Subsequently, in the inference phase, we implement the rendering process of Vosh based on WebGL, enabling real-time view synthesis on mobile devices. Next, we elaborate the algorithm details in the training and inference phases.
IV Hybrid Representation construction
The construction of a hybrid representation in the training phase commences with the initialization of a volume comprising solely of voxels. Then, by calculating the difference in rendering quality between the volume and its underlying surface, the density field corresponding to regions with similar rendering quality are meshed to replace the volumetric density values corresponding to the same positions.
IV-A Grid training for initial voxels
@@
Initially, the entire scene is reconstructed using a high-resolution voxel grid. It retains sufficient geometric and appearance information essential for subsequent mesh extraction and hybrid representation. Here, we employ SNeRG++ (an improved version of SNeRG [8] in MERF [18]) as the benchmark for our voxel grid, due to its capability to effectively model unbounded scenes and generate real-time rendering assets.
@@
Concretely, SNeRG++ is built upon the volume rendering principles of NeRF, which involves sampling along the direction of a ray originating from the ray origin $\mathbf{o}$ along the direction $\mathbf{d}$ to obtain sample points $\mathrm{\mathbf{x}}_{i}=\textbf{o}+t_{i}\mathrm{\mathbf{d}}$. Based on the spatial position of these sample points, it uses the neural radiance field to calculate volumetric density $\sigma_{i}$, resulting in rendering weights $w_{i}$:
$w_{i}=\alpha_{i}T_{i},\;\;\;\;\;\;\;\;T_{i}=\prod^{i-1}_{j=1}(1-\alpha_{j}),\;\;\;\;\;\;\;\;\alpha_{i}=1-e^{-\sigma_{i}\delta_{i}},$ (1)
@@
where $T_{i}$, $\alpha_{i}$ and $\delta_{i}$ represent the transmittance, opacity and spacing of sampling point $i$, respectively. SNeRG++ adopts the concept of deferred rendering, decomposing the appearance information of the neural radiance field into diffuse color $\mathrm{\mathbf{c}}_{d,i}$ and view-dependent features $\mathbf{f}_{i}$. By cumulatively sampling the appearance information along the ray, we can obtain the diffuse color $\mathbf{C}_{d}$ and view-dependent features $\mathbf{F}$ of the ray as
$\mathrm{\mathbf{C}}_{d}=\sum_{i}w_{i}\mathrm{\mathbf{c}}_{d,i},\;\;\;\;\;\;\;\;\mathrm{\mathbf{F}}=\sum_{i}w_{i}\mathrm{\mathbf{f}}_{i}.$ (2)
@@
Finally, the view-dependent color is calculated by inputting the diffuse color $\mathrm{\mathbf{C}}_{d}$, view-dependent features $\mathbf{F}$, and the ray direction $d$ into a miniature MLP. The result is then added to the diffuse color $\mathrm{\mathbf{C}}_{d}$ to obtain the final color $\mathbf{C}$ of the ray.
@@
Besides, in order to enhance the modeling of regions situated far from the center within unbounded scenes and facilitate efficient ray marching calculations, SNeRG++ incorporates a contraction function defined on piecewise projection.
IV-B Meshing density field
The motivation for meshing density field is to convert as many voxels as possible to mesh with minimal quality loss. Usually, there is a significant quality gap between voxel and mesh representation, but the distribution corresponding to the gap is unknown. We identify spatial regions where the volume and its implicit surface offer comparable rendering quality by measuring the quality difference between the surface after differentiable refinement and initial voxels, as well as meshing the density field within these regions. In this way, we can achieve speed gain with minimal quality loss.